# Loss Functions and Metrics in Deep Learning. A Review

### v1

@@ 7.1.8 Circle Loss
  • $s_{pos_{i}}$ and $s_{neg_{j}}$ represent the pairwise similarity between the positive and negative pairs. The positive pairs belong to the same class, and the negative pairs belong to different classes.
@@
  • $O_{pos_{i}}$ and $O_{neg_{j}}$ represent the user-defined margins for positive and negative pairs, respectively. $O_{pos}$ is a margin that should be smaller than the expected similarity of positive pairs, while $O_{neg}$ should be larger than the expected similarity of negative pairs. Thus, you want to choose $O_{pos}$ and $O_{neg}$ such that $O_{pos}<O_{neg}$.
  • $\alpha_{pos_{i}}$ and $\alpha_{neg_{j}}$ are slack variables that ensure the positive similarities are larger than $O_{pos_{i}}$ and the negative similarities are smaller than $O_{neg_{j}}$. This is achieved by setting the minimum value to 0, ignoring similarities that have already met the margin requirement.
  • $sum_{pos}$ and $sum_{neg}$ are exponentiated and scaled sums of the positive and negative similarities, respectively. The exponential function ensures that all values are positive and emphasize larger values, while $\gamma$ is a scaling factor that controls the rate at which the emphasis increases. The slack variables $\alpha_{pos_{i}}$ and $\alpha_{neg_{j}}$ are used in the exponent to give more weight to pairs far from meeting the margin requirement.
  • Finally, $L$ is the Circle Loss computed as the logarithm of 1 plus the product of $sum_{pos}$ and $sum_{neg}$. This encourages both $sum_{pos}$ and $sum_{neg}$ to be small, which in turn encourages the positive similarities to be large and the negative similarities to be small. Adding one inside the logarithm ensures that the argument is always positive, and the logarithm itself helps dampen the effect of large values and reduce the effect of outliers.
The circle loss has a more definite convergence target than the triplet loss because there is a single point in the ($s_{neg}$, $s_{pos}$) space toward which the optimization is driven ($O_{neg}$, $O_{pos}$). However, choosing ($O_{neg}$, $O_{pos}$) is arbitrary. In practice, it is common to use cross-validation or a separate validation set to tune these hyperparameters.
A common strategy is to start with small margins and gradually increase them. If the margins are too large, the model struggles to learn; if they are too small, it learns embeddings that are not discriminative enough. In some implementations of Circle Loss, $O_{pos}$ and $O_{neg}$ are not independent but related by a constant margin $m$, which is set to ensure a sufficient gap between positive and negative pairs. In this case, we only need to tune one of the margins or the gap $m$.
7.1.9 Barlow Twins Loss
@@
The Barlow Twins loss [168] is a self-supervised learning approach. The key idea is to make the outputs of a two-twin network, processing two different views of the same image, as similar as possible while reducing the redundancy between the dimensions of these representations. It encourages the network to learn highly informative features about the input and non-redundant.
@@
Given the batch size $N$ and the dimensionality $D$ of the embeddings, the network processes two different augmentations of the same input data, producing the embeddings $z_{a}$ and $z_{b}$. These embeddings are then normalized to have zero mean and unit variance along the batch dimension.
The computation of the Barlow Twins loss can be formulated as follows: 1. Compute the cross-correlation matrix $C$:
@@
2. Compute the difference matrix $C_{diff}$ between the cross-correlation matrix and an identity matrix and square it:
@@
3. Scale the off-diagonal elements of $C_{diff}$ by a factor $\lambda$:
@@
4. Finally, the Barlow Twins loss $L$ is the sum of all elements in the updated $C_{diff}$:
@@
By backpropagating this loss and updating the model’s parameters, the network is trained to minimize redundancy and increase the similarity between two different views of the same image, learning more robust and informative features as a result.
This method does not require a fixed number of classes and does not suffer from data expansion as it does not require explicit negative examples. However, the model in the paper required large dimensionality of the final representation for good performance, and the performance is not robust to remove certain distortions to the inputs [160].
7.1.10 SimSiam Loss
SimSiam [169] is a self-supervised learning aimed at learning representations by pushing two views of the same image to be as similar as possible. While it is not explicitly designed for face recognition, it can be used to learn meaningful representations of faces.
Given two different augmentations $x_{1}$ and $x_{2}$ of the same image, they are forwarded through a neural network and a prediction head to obtain the features $z_{1},z_{2}$ and the predictions $p_{1},p_{2}$.
The loss is defined as the negative cosine similarity between the predictions and the features of the different augmentations.
$L=-\frac{1}{2}[\frac{(p_{1}^{T}z_{2})}{(||p_{1}||\cdot||z_{2}||)}+\frac{(p_{2}^{T}z_{1})}{(||p_{2}||\cdot||z_{1}||)}],$ (69)
where ${}^{T}$ denotes transpose, and |||||$\cdot$||| | ⋅ | | denotes the L2 norm.
This loss encourages the model to make the predictions $p_{1}$ and $p_{2}$ as similar as possible to the features $z_{2}$ and $z_{1}$, respectively.
An important part of the SimSiam approach is a stop-gradient operation applied to $z_{2}$ in the first term and $z_{1}$ in the second term of the loss function. This means that the gradients are not backpropagated through these variables during training. The stop-gradient operation is critical to avoid the model collapsing into trivial solutions where the features and the predictions are the same.
The advantages of the SimSiam loss are:
@@ 7.1.10 SimSiam Loss
    No Negative Pair: Unlike contrastive learning methods that require negative examples, SimSiam does not require any. This simplifies the model training process and can make it more efficient.
@@
    Stop-Gradient Operation: The use of stop-gradient operation in SimSiam makes it less prone to collapsing into trivial solutions. This is a significant advantage because many self-supervised learning models struggle with this problem.
@@
    Simplicity: SimSiam is simpler compared to other self-supervised learning methods. It uses a symmetric architecture and a simple loss function which encourages two views of the same image to have similar representations.
@@
    Hyperparameter Sensitivity: SimSiam has a few crucial hyperparameters, such as the learning rate and weight decay, which require careful tuning to get the best performance. Incorrect settings can significantly degrade the model’s performance.
@@
    Dependence on Data Augmentation: The success of SimSiam, like many other self-supervised learning models, heavily relies on the choice and quality of data augmentations. This requires domain knowledge and potentially significant effort to determine the most effective augmentations.
@@
    Non-semantic Features: One common issue with self-supervised learning methods, including SimSiam is that the features learned may not necessarily be semantically meaningful. They are good at capturing low-level features but may not be as effective in capturing high-level semantic information.
8 Image Generation
Image generation in deep learning involves using artificial neural networks to generate new images. This task has been revolutionized by developing various models such as Variational Autoencoders (VAEs) [170, 171, 172, 173], Generative Adversarial Networks (GANs) [174, 175, 176, 177, 178], Normalized Flow models (NFs) [179, 180, 181, 182, 183], Energy-Based Models (EBMs) [184, 185, 186, 187, 188, 189], and Diffusion Models [190, 191, 192, 193, 194]. These models allow the generation of high-quality images that can be used in various applications such as image super-resolution [195, 196, 197, 198], denoising [199, 200, 201], inpainting [202, 203, 204, 205], and style transfer [206, 207, 208, 209].
Variational Autoencoders (VAEs) are generative models that use deep learning techniques to create new data and learn latent representations of the input data. VAEs consist of two primary components: an encoder and a decoder. The encoder takes input data and compresses it into a lower-dimensional latent space, capturing the essential features of the data. This is typically a probabilistic process, producing a mean and standard deviation representing potential latent space values distribution. The decoder then takes a point from this latent space distribution and reconstructs the original data. The entire process is trained in such a way as to minimize the difference between the original and the reconstructed data, as well as to ensure the latent space approximates a standard Gaussian distribution. VAEs can generate new data by feeding the decoder points sampled from the latent space.
@@ 8 Image Generation
Normalizing Flow Models (NFs) can construct complex probability distributions by transforming a simple base distribution through a sequence of invertible and differentiable transformations, or flows. These flows warp and twist the data space, allowing the model to generate diverse outputs. The parameters of these transformations are learned from data using maximum likelihood estimation. An advantage of Normalizing Flows over other generative models is their ability to provide an exact and tractable likelihood for a given sample, enabling efficient sampling and density estimation. However, they can be computationally intensive to train due to the need to compute and backpropagate through the Jacobian of the flows.
Energy-Based Models (EBMs) learn a scalar energy function to distinguish real data points from unlikely ones, assigning lower energy values to points similar to the training data and higher values otherwise. A neural network often parameterizes this function and learns from the data. Sampling in EBMs is typically done via Markov Chain Monte Carlo (MCMC) [211] methods, producing samples distributed according to the learned energy function. While EBMs can represent a wide variety of data distributions, they can be challenging to train due to the intractability of the partition function and the computational expense of MCMC sampling.
Diffusion models can create new data by simulating a random process, specifically a diffusion process. The process starts with a simple data distribution (e.g., Gaussian noise) and gradually transforms it towards the target data distribution, like a particle undergoing diffusion or a random walk. This transformation is controlled by a neural network trained to model the reverse process, taking real-world data and applying a series of transformations to reduce it to noise. During generation, an initial noise sample is iteratively updated using this reverse process model, running the process backward, leading to a sample from the target data distribution. This approach allows the generation of complex and high-quality data, such as images, by simulating a smooth transition from noise to the desired data.
@@
8.1 Image Generation Loss functions
The loss function in a VAE consists of the reconstruction loss and the Kullback-Leibler (KL) (refer to section 8.1.2). The reconstruction loss measures how well the decoded data matches the original input data. The KL divergence measures how much the learned distribution in the latent space deviates from a target distribution, usually a standard normal distribution. KL-divergence is used as a regularization term to ensure that the distributions produced by the encoder remain close to a unit Gaussian, penalizing the model if the learned distributions depart from it.
The most common loss function used in GANs is the adversarial loss, which is the sum of the cross-entropy loss between the generator’s predictions and the real or fake labels. More recently, WGAN [212] applied the Wasserstein distance as an alternative to training GANs to improve the stability and avoid mode collapse that occurs when the generator network stops learning the underlying data distribution and begins to produce a limited variety of outputs, rather than a diverse range of outputs that accurately represent the true data distribution.
@@ 8.1 Image Generation Loss functions
8.1.1 Reconstruction Loss
The purpose of the reconstruction loss is to ensure that the decoder network of the VAE can reconstruct an input image that is as close as possible to the original image. In essence, the reconstruction loss compares the original image to the reconstructed image and penalizes any difference between the two. The reconstruction loss is calculated as the mean squared error (MSE) between the original image and the reconstructed image, which is defined as follows
@@ 8.1.1 Reconstruction Loss
where $x_{i}$ is the original image, $\hat{x_{i}}$ is the reconstructed image, and $N$ is the number of images.
Another metric that can be used for reconstruction loss is binary cross-entropy (BCE), a common choice for binary images. The BCE loss can be defined as:
$L_{BCE}=-\sum_{i=1}^{n}y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}}),$ (71)
where $y_{i}$ is the binary value of the original image and $\hat{y_{i}}$ is the binary value of the generated image.
In the context of deep learning, reconstruction loss has been used in various image generation tasks such as image restoration, image generation, and image synthesis.
8.1.2 Kullback-Leibler Divergence Loss
@@
8.1.3 Adversarial Loss
The adversarial loss is the main function used in Generative Adversarial Networks (GANs). It is based on the idea of a minimax game between the two neural networks. The adversarial loss is used to train the generator, defined as the negative log-likelihood of the discriminator’s output.
Goodfellow et al. first introduced the adversarial loss in [174]. The authors showed that this loss function leads to the convergence of the generator to a Nash equilibrium, where the generator produces realistic images that the discriminator cannot distinguish from real images.
The adversarial loss can be defined as:
$L_{adv}(G,D)=-\mathbb{E}{x\sim p{data}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$ (73)
where $G$ is the generator network, $D$ is the discriminator network, $x$ are real images from the data distribution $p_{data}(x)$, $z$ are random noise inputs from the noise distribution $p_{z}(z)$, and $G(z)$ are generated images.
The adversarial loss encourages the generator to generate indistinguishable images from real images and the discriminator to correctly classify real images and generated images. The training process alternates between updating the generator and the discriminator until the generator produces realistic images.
8.1.4 Wasserstein loss
The Wasserstein loss [212] is used in generative models, such as Generative Adversarial Networks (GANs), to measure the difference between the real and generated distributions. The Wasserstein loss is the minimum work required to transform one probability distribution into another. The amount of work is the sum of the distances between each sample multiplied by their probability mass. The Wasserstein loss can be written as
@@ 8.1.4 Wasserstein loss
where $p_{data}$ and $p_{gen}$ represent the real and generated distributions, respectively. $\Gamma(p_{data},p_{gen})$ is the set of joint distributions with marginals $p_{data}$ and $p_{gen}$, and $\gamma(x,y)$ is the joint distribution of $(x,y)$.
The Wasserstein loss has been widely used in GANs for its ability to provide a more stable training process and to avoid the mode collapse problem, where the generator produces a limited number of similar outputs.
8.1.5 Negative Log-likelihood in Normalizing Flows
@@
In Normalizing Flows, the objective is to learn a transformation (or a sequence of transformations) that maps a complex data distribution to a simpler distribution, for example, a standard Gaussian. The transformations are chosen such that their Jacobian determinant is easy to compute, which allows using the change of variables formula to compute the likelihood of the data under the model.
@@
Given a data point $x$, let’s denote $z=f_{\theta}(x)$ as the mapping of $x$ to the base distribution under the flows parameterized by $\theta$, and $p_{z}(z)$ as the density of the base distribution. Then, the log-likelihood of $x$ under the model is given by
$\log p_{\theta}(x)=\log p_{z}(f_{\theta}(x))-\log\left|\det\frac{\partial f_{\theta}(x)}{\partial x}\right|$ (75)
The second term is the log absolute determinant of the Jacobian of the transformation, which corrects for the change in volume due to the transformation.
@@
The loss function that we minimize during training is the negative log-likelihood of the data. If our dataset consists of $N$ points $x_{1},x_{2},\ldots,x_{N}$, then the loss function $L(\theta)$ is expressed as
$L(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log p_{\theta}(x_{i})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log p_{z}(f_{\theta}(x_{i}))-\log\left|\det\frac{\partial f_{\theta}(x_{i})}{\partial x_{i}}\right|\right]$ (76)
In practice, stochastic gradient descent or a variant minimizes this loss function and learns the parameters $\theta$ of the flows.
Normalizing Flows are computationally intensive to train, due to the need to compute and backpropagate through the Jacobian of the flows. For this reason, methods such as RealNVP [179] and Glow [180] are designed so that the determinant of the Jacobian can be computed efficiently.
8.1.6 Contrastive Divergence
Contrastive Divergence (CD) can be used in training Energy-Based Models (EBMs), specifically for estimating the log-likelihood gradient. As with RBMs, the objective is to maximize the likelihood of the data under the model. The key difference is that the energy function is defined over the visible and hidden units in EBMs, whereas it’s defined over the joint configuration of visible and hidden units in RBMs.

### v3

@@ 5. The final A-Softmax loss $L$ is calculated as the negative log-likelihood of the correct class:
  • $s_{pos_{i}}$ and $s_{neg_{j}}$ represent the pairwise similarity scores for positive pairs (same class) and negative pairs (different classes), respectively.
@@
  • $O_{pos}$ and $O_{neg}$ are user-defined margins for positive and negative pairs. Typically, $O_{pos}$ is set smaller than the expected similarity for positive pairs, and $O_{neg}$ is set larger than the expected similarity for negative pairs, ensuring $O_{pos}<O_{neg}$.
  • $\alpha_{pos_{i}}$ and $\alpha_{neg_{j}}$ are slack variables ensuring that the similarity scores for positive pairs are greater than $O_{pos}$, and for negative pairs are less than $O_{neg}$. These are computed using $\max(0,\cdot)$ to ignore pairs that already satisfy the margin requirements.
  • $\text{sum}_{pos}$ and $\text{sum}_{neg}$ are the scaled and exponentiated sums of positive and negative similarities, respectively. The exponential scaling, modulated by $\gamma$, emphasizes differences, particularly for pairs that do not meet the margin requirement. $\gamma$ acts as a scaling factor controlling this emphasis.
  • The overall loss $L$ is calculated as the logarithm of $1+\text{sum}_{pos}\cdot\text{sum}_{neg}$. This formulation drives $\text{sum}_{pos}$ to be low (favoring high similarity for positive pairs) and $\text{sum}_{neg}$ to be low (favoring low similarity for negative pairs). The logarithmic transformation stabilizes the loss, mitigating the impact of outliers and ensuring the argument remains positive.
Circle loss offers a clear convergence target in the similarity space defined by ($s_{neg}$, $s_{pos}$), directing optimization toward the point ($O_{neg}$, $O_{pos}$). However, selecting the appropriate margins ($O_{neg}$, $O_{pos}$) can be challenging and typically requires hyperparameter tuning, often through cross-validation. A common practice is to begin with smaller margins and gradually increase them, avoiding overly large margins that can hinder learning or excessively small margins that fail to enforce sufficient discrimination.
In some implementations, $O_{pos}$ and $O_{neg}$ are not independent but are related by a constant margin $m$, simplifying the tuning process. By setting a fixed gap between positive and negative pairs, this approach ensures a consistent separation, making it necessary to only tune one of the margins or the margin gap $m$.
Barlow Twins Loss
@@
The Barlow Twins loss [196] is a self-supervised learning approach designed to learn highly informative and non-redundant representations. The main idea is to make the outputs of a twin network, processing two different augmentations of the same image, as similar as possible while reducing redundancy between the dimensions of these representations.
@@
Given a batch size $N$ and the dimensionality $D$ of the embeddings, the network processes two different augmentations of the same input data, producing the embeddings $z_{a}$ and $z_{b}$. These embeddings are then normalized to have zero mean and unit variance along the batch dimension, resulting in $z_{a_{norm}}$ and $z_{b_{norm}}$.
The computation of the Barlow Twins loss involves the following steps:
1. Compute the Cross-Correlation Matrix $C$:
@@
2. Compute the Difference Matrix $C_{diff}$: Subtract the identity matrix $I$ from $C$ and square the elements:
@@
3. Scale Off-Diagonal Elements: Apply a scaling factor $\lambda$ to the off-diagonal elements of $C_{diff}$:
@@
4. Compute the Barlow Twins Loss $L$: Sum all the elements in the adjusted $C_{diff}$:
@@
The loss function is designed to achieve two main goals: (1) making the diagonal elements of the cross-correlation matrix close to 1, which ensures that the representations of the two views are highly correlated; and (2) making the off-diagonal elements close to 0, which reduces redundancy among different components of the learned representations.
This method does not rely on a fixed number of classes and avoids data expansion issues as it does not require explicit negative examples. However, achieving good performance often requires a large dimensionality of the final representation. Additionally, the model’s robustness can be affected by certain distortions to the inputs, as noted in [189].
Barlow Twins loss is effective in scenarios where it is important to learn representations that are both invariant to different augmentations of the same input and decorrelated across different dimensions, making it suitable for various self-supervised learning tasks.
SimSiam Loss
SimSiam [197] is a self-supervised learning method designed to learn meaningful representations by encouraging the similarity between two augmented views of the same image. Although not explicitly designed for face recognition, SimSiam can be employed to learn robust face representations.
Given two different augmentations, $x_{1}$ and $x_{2}$, of the same image, these are passed through a neural network and a prediction head to obtain feature embeddings $z_{1},z_{2}$ and predictions $p_{1},p_{2}$. The loss function is defined as the negative cosine similarity between the predictions and the corresponding features from the alternate view:
$L=-\frac{1}{2}\left[\frac{p_{1}^{T}z_{2}}{||p_{1}||\cdot||z_{2}||}+\frac{p_{2}^{T}z_{1}}{||p_{2}||\cdot||z_{1}||}\right],$ (91)
where T denotes the transpose, and |||||$\cdot$||| | ⋅ | | denotes the L2 norm.
This loss encourages the model to align the predictions $p_{1}$ and $p_{2}$ with the features $z_{2}$ and $z_{1}$, respectively, from different augmented views. An essential aspect of SimSiam is the stop-gradient operation applied to $z_{2}$ in the first term and $z_{1}$ in the second term of the loss function. This operation prevents gradients from backpropagating through these feature vectors, which is crucial for avoiding trivial solutions where the model could simply collapse by producing identical outputs for all inputs.
The advantages of SimSiam Loss are the following:
@@
    No Need for Negative Pairs: Unlike contrastive learning methods that require negative examples, SimSiam does not need such pairs, simplifying the training process and potentially improving efficiency.
@@
    Stop-Gradient Operation: The stop-gradient operation is a key feature that helps prevent the collapse into trivial solutions, a common problem in self-supervised learning.
@@
    Simplicity: SimSiam employs a straightforward loss function and a symmetric network architecture, making it simpler compared to other self-supervised methods. It focuses on making representations from different views of the same image similar.
@@
    Hyperparameter Sensitivity: The method’s performance is sensitive to hyperparameters like learning rate and weight decay, necessitating careful tuning to achieve optimal results. Incorrect settings can significantly impact performance.
@@
    Dependence on Data Augmentation: The effectiveness of SimSiam heavily relies on the quality and choice of data augmentations, which requires domain expertise and considerable effort to optimize.
@@
    Non-semantic Features: A limitation common to many self-supervised learning methods, including SimSiam, is that the learned features may not always capture high-level semantic information. They might excel in identifying low-level patterns but may lack in semantic depth.
SimSiam offers an approach to self-supervised learning by leveraging the similarity between augmented views without the need for negative samples, but it also requires careful consideration of augmentations and hyperparameters to be most effective.
4.5 Image Generation
Image generation in deep learning involves using artificial neural networks to generate new images. This task has progressed significantly with models such as Variational Autoencoders (VAEs) [198, 199, 200, 201], Generative Adversarial Networks (GANs) [202, 203, 204, 205, 206], Normalized Flow models (NFs) [207, 208, 209, 210, 211], Energy-Based Models (EBMs) [212, 213, 214, 215, 216, 217], and Diffusion Models [218, 219, 220, 221, 222]. These models can generate high-quality images that can be used in various applications such as image super-resolution [223, 224, 225, 226], denoising [227, 228, 229], inpainting [230, 231, 232, 233], and style transfer [234, 235, 236, 237].
Variational Autoencoders (VAEs) are generative models that use deep learning techniques to create new data and learn latent representations of the input data. They consist of an encoder and a decoder. The encoder compresses input data into a lower-dimensional latent space, while the decoder reconstructs the original data from points in the latent space. The process is trained to minimize the difference between original and reconstructed data and ensure the latent space approximates a standard Gaussian distribution. VAEs can generate new data by feeding the decoder points sampled from the latent space.
@@ 4.5 Image Generation
Normalizing Flows use invertible transformations to generate diverse outputs and provide an exact and tractable likelihood for a given sample, enabling efficient sampling and density estimation. However, they can be computationally intensive to train.
Energy-Based Models (EBMs) learn a scalar energy function to distinguish real data points from unlikely ones. A neural network often parameterizes this function and learns from the data. Sampling in EBMs is typically done via Markov Chain Monte Carlo (MCMC) methods. EBMs can represent a wide variety of data distributions but can be challenging to train due to the intractability of the partition function and the computational expense of MCMC sampling.
Diffusion models use a random process to transform simple data distribution, like Gaussian noise, into the desired target data distribution. This is controlled by a trained neural network, allowing the generation of high-quality data, like images, through a smooth transition from noise to the desired data.
@@
4.5.1 Image Generation Loss functions
The loss function in a VAE consists of the reconstruction loss and the Kullback-Leibler Divergence Loss (KL) (refer to section 4.5.1). The reconstruction loss measures how well the decoded data matches the original input data. The KL divergence measures how much the learned distribution in the latent space deviates from a target distribution, usually a standard normal distribution. KL-divergence is used as a regularization term to ensure that the distributions produced by the encoder remain close to a unit Gaussian, penalizing the model if the learned distributions depart from it.
The most common loss function used in GANs is the adversarial loss, which is the sum of the cross-entropy loss between the generator’s predictions and the real or fake labels. Later, WGAN [239] applied the Wasserstein distance as an alternative to training GANs to improve stability and avoid mode collapse that occurs when the generator network stops learning the underlying data distribution and begins to produce a limited variety of outputs, rather than a diverse range of outputs that accurately represent the true data distribution.
@@ 4.5.1 Image Generation Loss functions
Reconstruction Loss
Reconstruction loss is a essential in tasks involving image generation and reconstruction, such as in Variational Autoencoders (VAEs), image restoration, and image synthesis. The primary purpose of the reconstruction loss is to ensure that the output from the decoder network is as close as possible to the original input image. This loss penalizes the differences between the original and reconstructed images, encouraging the model to learn a faithful representation of the input data.
One common form of reconstruction loss is the Mean Squared Error (MSE), which measures the average squared differences between the original and reconstructed images. The MSE loss is defined as:
@@
where $x_{i}$ is the original image, $\hat{x_{i}}$ is the reconstructed image, and $N$ is the number of images in the batch. The MSE loss is particularly suitable for continuous-valued image data, as it directly quantifies the pixel-wise differences.
For binary images, Binary Cross-Entropy (BCE) loss is often used, as it is more appropriate for modeling binary outputs. The BCE loss can be defined as:
$L_{BCE}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right],$ (93)
where $y_{i}$ is the binary value of the original image, and $\hat{y_{i}}$ is the binary value of the reconstructed image. The BCE loss measures the divergence between the true binary values and the predicted probabilities, making it suitable for tasks involving binary or probabilistic outputs.
Reconstruction loss is widely used in various deep learning tasks beyond VAEs, including:
  1. 1.
    Image Restoration: In tasks such as denoising, super-resolution, and inpainting, reconstruction loss helps the model generate cleaner, high-resolution, or complete versions of images.
  2. 2.
    Image Generation: Generative models like GANs and VAEs use reconstruction loss to ensure that generated images resemble real samples.
  3. 3.
    Anomaly Detection: Reconstruction loss is used in models designed to detect anomalies by measuring how well the model reconstructs normal versus anomalous data.
Kullback-Leibler Divergence Loss
@@
Adversarial Loss
Adversarial loss is the fundamental loss function used in Generative Adversarial Networks (GANs). It is derived from a minimax game framework involving two neural networks: the generator ($G$) and the discriminator ($D$). The adversarial loss is essential for training the generator, as it aims to generate data indistinguishable from real data, while the discriminator learns to differentiate between real and generated data.
The concept of adversarial loss was first introduced by Goodfellow et al. in [202]. The authors demonstrated that this loss function facilitates the convergence of the generator towards a Nash equilibrium, where the generator’s output is so realistic that the discriminator cannot reliably distinguish between real and generated data.
The adversarial loss for a GAN can be expressed as:
$L_{adv}(G,D)=-\mathbb{E}_{x\sim p_{data}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))],$ (95)
where:
  • $G$ represents the generator network that maps a noise vector $z$ to a data space, generating an image $G(z)$.
  • $D$ represents the discriminator network that outputs the probability that an input image is real.
  • $x$ denotes real images drawn from the real data distribution $p_{data}(x)$.
  • $z$ denotes latent vectors sampled from a noise distribution $p_{z}(z)$.
In this setup, the generator aims to minimize the adversarial loss by producing realistic images $G(z)$ that fool the discriminator, while the discriminator tries to maximize the loss by correctly identifying real images from generated ones. This adversarial objective can be framed as a two-player game with the following optimization goals:
For the Discriminator $D$, maximize the probability of correctly classifying real and generated samples:
$\max_{D}\mathbb{E}_{x\sim p_{data}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (96)
For the Generator $G$, minimize the likelihood of the discriminator identifying the generated samples as fake:
$\min_{G}\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (97)
In practice, the generator is often trained to maximize $\mathbb{E}_{z\sim p_{z}(z)}[\log D(G(z))]$, which can help mitigate issues such as vanishing gradients when the discriminator becomes too strong.
The training process involves alternately updating the generator and the discriminator. The generator improves its output to produce more realistic images, while the discriminator becomes better at distinguishing real from generated data. This dynamic continues until the generator’s images are indistinguishable from real ones, ideally reaching an equilibrium where the discriminator can no longer differentiate between real and fake samples with certainty.
Adversarial loss is an effective tool for training GANs, enabling the generation of highly realistic data. However, it can also lead to challenges such as mode collapse and instability during training, requiring careful balancing and regularization techniques.
Wasserstein Loss
The Wasserstein loss, introduced by Arjovsky et al. in [239], is used in generative models like Generative Adversarial Networks (GANs) to measure the divergence between the real data distribution and the generated data distribution. Unlike the traditional adversarial loss, Wasserstein loss is based on the Wasserstein distance (also known as Earth Mover’s Distance), which quantifies the cost of transporting mass from one distribution to match another.
Mathematically, the Wasserstein loss is defined as:
@@
where:
  • $p_{data}$ and $p_{gen}$ represent the real and generated data distributions, respectively.
  • $\Gamma(p_{data},p_{gen})$ is the set of all joint distributions $\gamma(x,y)$ whose marginals are $p_{data}$ and $p_{gen}$.
  • $\gamma(x,y)$ represents a coupling between $p_{data}$ and $p_{gen}$, effectively describing how mass should be transported to transform one distribution into the other.
  • $||x-y||$ is a chosen metric (typically Euclidean distance) between samples $x$ and $y$ from the real and generated distributions, respectively.
The Wasserstein distance measures the minimum work required to transform the generated distribution into the real distribution, where work is defined as the integral of the distance between points $x$ and $y$, weighted by the joint distribution $\gamma(x,y)$. This distance provides a meaningful measure even when the real and generated distributions have disjoint supports, which is a common issue with traditional GAN losses.
The Wasserstein loss has a couple of advantages over the adversarial loss. On the one hand, the Wasserstein loss tends to produce more stable training dynamics compared to traditional adversarial losses. This stability arises because the Wasserstein distance provides continuous and differentiable gradients, even when the support of the distributions does not overlap. On the other hand, unlike traditional GANs, which can suffer from mode collapse (where the generator produces a limited variety of outputs), Wasserstein GANs (WGANs) mitigate this problem. The Wasserstein distance ensures that the generator is penalized proportionally to the distance between distributions, encouraging the generation of diverse outputs.
The implementation of Wasserstein loss in GANs typically requires some modifications, such as using weight clipping or gradient penalty (as seen in WGAN-GP [244]) to enforce the Lipschitz constraint on the discriminator’s function. These modifications are necessary to ensure the validity of the distance measure and the stability of the training process.
Negative Log-Likelihood in Normalizing Flows
@@
Normalizing Flows are a class of generative models that aim to learn a bijective transformation (or a series of transformations) that maps a complex data distribution to a simpler, typically standard, base distribution, such as a Gaussian. The key feature of Normalizing Flows is the ability to compute the exact likelihood of the data by leveraging the change of variables formula, which requires the transformation to be invertible and differentiable with a tractable Jacobian determinant.
@@
Given a data point $x$, let $z=f_{\theta}(x)$ denote the transformation of $x$ to the latent space under the flow model parameterized by $\theta$, and $p_{z}(z)$ denote the density of the base distribution. The log-likelihood of $x$ under the model is computed as:
$\log p_{\theta}(x)=\log p_{z}(f_{\theta}(x))+\log\left|\det\frac{\partial f_{\theta}(x)}{\partial x}\right|,$ (99)
where the second term, $\log\left|\det\frac{\partial f_{\theta}(x)}{\partial x}\right|$, is the log absolute determinant of the Jacobian matrix of the transformation $f_{\theta}$. This term accounts for the change in volume induced by the transformation, ensuring the correct density transformation from the base distribution to the data distribution.
@@
The training objective for Normalizing Flows involves maximizing the likelihood of the data, which is equivalent to minimizing the negative log-likelihood. Given a dataset with $N$ data points $x_{1},x_{2},\ldots,x_{N}$, the loss function $L(\theta)$ is expressed as:
$L(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log p_{\theta}(x_{i})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log p_{z}(f_{\theta}(x_{i}))+\log\left|\det\frac{\partial f_{\theta}(x_{i})}{\partial x_{i}}\right|\right].$ (100)
During training, this loss function is minimized using stochastic gradient descent or its variants, allowing the model to learn the parameters $\theta$ of the flow transformations.
Normalizing Flows can be computationally intensive to train due to the necessity of computing and backpropagating through the Jacobian of the transformations. Consequently, specialized architectures such as RealNVP [207] and Glow [208] are designed to facilitate efficient computation of the Jacobian determinant, often using techniques like affine coupling layers and invertible $1\times 1$ convolutions, which simplify the computation of the determinant and its gradient.
Contrastive Divergence